A fleet of delivery vans will cost $160,000. Depreciation is straight line over five years. For capital budgeting purposes, however, you will use cash flow from salvage at the end of four years and assuming you can obtain a sales price of one half of the initial cost. What will that Year 4 savage cash flow be if the company's tax rate is 27.0%?
Answer #1

Written down value after 4 years = Initial Cost - Depreciation for 4 years

= 160,000 - 160,000*4/5

= $32,000

Selling price = $80,000

Gain on Sale = $48,000

Tax on gain = 48000*27% = $12,960

Salvage cash flow = Sales value - tax on gain

= 80,000 - 12,960

= $67,040